OpenStudy (anonymous):

Can you factor this? i'm getting confused if i can for not. (x-x^-2)

OpenStudy (anonymous):

\[x-x^2\] both terms have a common factor of \(x\)

OpenStudy (anonymous):

no its raised to the -2

OpenStudy (anonymous):

\[x-x ^{-2}\]

OpenStudy (anonymous):

use mathway.com

OpenStudy (anonymous):

\[x-\frac{1}{x^2}\] is what you have in that case

OpenStudy (anonymous):

you can write it as a fraction if you like and get \[\frac{x^3-1}{x^2}\] or even \[\frac{(x-1)(x^2+x+1)}{x^2}\] if that helps any not sure what the goal of factoring this is

OpenStudy (whpalmer4):

@satellite73 user might just be trying to figure out if something is fully factored or not...